Consider the following reaction where Kc = 6.50×10-3 at 298 K: 2 NOBr (g) goes to 2 NO (g) + Br2 (g) A reaction mixture was found to contain 8.54×10-2 moles of NOBr (g), 2.30×10-2 moles of NO (g), and 4.35×10-2 moles of Br2 (g), in a 1.00 liter container. Indicate True (T) or False (F) for each of the following:

1. In order to reach equilibrium NOBr(g) must be consumed .

2. In order to reach equilibrium Kc must increase .

3. In order to reach equilibrium NO must be consumed .

4. Qc is greater than Kc.

5. The reaction is at equilibrium. No further reaction will occur.
Answer #1

Qc = [NO]^2[Br2] / [NOBr]^2
= (2.30*10^-2)^2 * (4.35*10^-2) / (8.54*10^-2)^2
= (2.30*10^-5) / (7.29*10^-3)
= 3.16*10^-3

This is less than Kc.
So, reaction will move in forward direction to achieve equilibrium.

So, concentration of NO and Br2 will increase and that of NOBr will decrease

Answer:
1: T
2: F (Kc Will not change since temperature is not changing)
3: F
4: F
5: F